Consider a equation for t below 191×2=3.97t×t+33.4t
What value of t satisfies this equation
1
Expert's answer
2020-01-20T05:24:03-0500

Given that,

"\Rightarrow 191×2=3.97t×t+33.4t"

"\Rightarrow 382=3.97t^2+33.4t"

"\Rightarrow 3.97t^2+33.4t-382=0"

it is the quadratic equation, so it will have 2 roots.

Now, from the quadratic equation formula,


"\Rightarrow t=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}"


"\Rightarrow t=\dfrac{-33.4\pm\sqrt{(33.4)^2+4\times 3.97\times382}}{2\times 3.97}"


"\Rightarrow t=\dfrac{-33.4\pm\sqrt{1115.56+6066.16}}{7.94}"


"\Rightarrow t=\dfrac{-33.4\pm \sqrt{7181.72}}{7.94}"


"\Rightarrow t=\dfrac{-33.4\pm84.74}{7.94}"

Taking "+",


"\Rightarrow t=\dfrac{-33.4+84.74}{7.94}=6.46"

so,

"t=6.5"

"\u200b\u200b"

Taking "-"

"t=\dfrac{-118.14}{7.94}=-14.88"